# Stability Analysis of Geometric Control for a Canonical Class of Underactuated Aerial Vehicles with Spurious Forces

### v1

@@ I Introduction
Distinct from platform-specific solutions, the control framework proposed by Michieletto et al. [14] addresses a significantly broader class of systems, referred to as generically tilted multi-rotors, admitting an arbitrary number of actuators and diverse geometric configurations. The objective of this line of work is to enable static hovering even in the presence of propeller failures, including configurations in which translational and rotational dynamics are intrinsically coupled. The proposed approach departs from the classical assumption that the control force is orthogonal to the plane containing the propeller centers and that control moments can be generated independently of the thrust force. Instead, the authors consider the more general case in which the direction of the control force is not fixed and control moments may induce translational effects. Within this framework, they introduce the concept of zero-moment direction (also referred to as preferential direction), defined as a virtual direction along which the intensity of the control force can be freely assigned while maintaining a zero net control moment. While this formulation provides a potentially unifying control paradigm for a wide range of multirotor configurations, its stability properties are supported primarily by heuristic arguments. Nevertheless, extensive experimental validation has demonstrated a remarkably large basin of attraction in fault scenarios. In particular, the controller was shown to successfully recover stability despite transient attitude errors approaching $50^{\circ}$, angular rates exceeding $50^{\circ}/\mathrm{s}$, and velocity errors above $1\,\mathrm{m/s}$ [15]. Despite this strong empirical evidence, a formal and rigorous stability analysis for this broad class of floating rigid bodies is still missing.
A subsequent work [13] advances a theoretical characterization of the above problem by providing sufficient conditions for local asymptotic stability of a hovering equilibrium within the same general modeling framework. The analysis, however, is developed under feasibility requirements on the actuation capabilities that effectively restrict the result to configurations in which the translational–rotational coupling does not yield intrinsically critical effects in the translational dynamics. In particular, the underlying assumptions ensure that the translational component induced by moment allocation can be locally eliminated by an equivalent input parametrization. Accordingly, the stability guarantees apply to platforms that can be treated as effectively free of spurious forces, whereas a rigorous analysis for configurations in which such forces are unavoidable and dynamically relevant remains open.
More generally, coupling between translational and rotational dynamics should be regarded as a system-level phenomenon that may, in principle, affect a broad class of floating vehicles. Table I summarizes the existing literature on floating aerial vehicles affected by force–moment coupling, highlighting the scope and limitations of current approaches. Early contributions addressed this coupling through approximate or dynamic feedback linearization, where the coupling terms are explicitly retained in the model but mitigated via dynamic extensions of the control input, as originally proposed for VTOL aircraft and helicopter systems by Hauser and co-authors and by Koo and Sastry [3, 7]. More recent works have instead adopted geometric control formulations that explicitly accommodate force–moment coupling. Among them, Zhong et al. [19] established almost global exponential stability for trajectory tracking of quadrotors with tilted propellers, covering configurations with force allocation matrices of rank one, two, and three. However, the resulting analysis remains inherently restricted to four-rotor platforms and does not naturally extend to vehicles with different actuator topologies or to more general classes of floating rigid bodies exhibiting structural force–moment coupling.
@@
The above discussion highlights the lack of topology-independent and rigorously established stability results for hovering control in the presence of intrinsic force–moment coupling.
@@
In this work, we embrace this viewpoint by considering the hovering stabilization problem for a canonical model that abstracts away the specific actuator topology, assuming only the existence of a structural coupling between control moments and translational forces. Our contribution provides, for the first time, a rigorous Lyapunov-based stability proof for the geometric control law proposed by [14], extending its validity to the abstract canonical model presented in Section II. In contrast to [13], our analysis relaxes the feasibility requirements underlying their stability proof, thereby covering a broader class of floating vehicles while remaining independent of the specific actuator topology. This result closes the gap between prior empirical evidence and theoretical certification. We prove that the controller guarantees local exponential stability of the hovering equilibrium. This generality comes at the cost of focusing on static hovering equilibria and does not address time-varying trajectory tracking.
@@
While the proposed analysis is inspired by the seminal work of Lee et al. [8, 9], the presence of spurious forces fundamentally alters the system structure. The induced non-minimum-phase behavior and intrinsic coupling violate the skew-symmetry and cascade properties on which the standard analysis relies. Consequently, we introduce a reformulated Lyapunov argument with new bounds specifically designed to handle the additional coupling terms. This work focuses exclusively on these theoretical properties, serving as the formal counterpart to previous experimental studies.
@@
The remainder of the work is organized as follows. Section II introduces the generic canonical model and summarizes the control law. Section III details the stability analysis. Concluding remarks are provided in Section IV.

### v2

@@ I Introduction
Table I summarizes the existing literature on floating aerial vehicles affected by force–moment coupling, highlighting the scope and limitations of current approaches. Early contributions addressed this coupling through approximate or dynamic feedback linearization, where the coupling terms are explicitly retained in the model but mitigated via dynamic extensions of the control input, as originally proposed for VTOL aircraft and helicopter systems by Hauser and co-authors and by Koo and Sastry [3, 7]. More recent works have instead adopted geometric control formulations that explicitly accommodate force–moment coupling. Among them, Zhong et al. [19] established almost global exponential stability for trajectory tracking of quadrotors with tilted propellers, covering configurations with force allocation matrices of rank one, two, and three. However, the resulting analysis remains inherently restricted to four-rotor platforms and does not naturally extend to vehicles with different actuator topologies or to more general classes of floating rigid bodies exhibiting structural force–moment coupling.
@@
Distinct from platform-specific solutions, the control framework proposed by Michieletto et al. [14] addresses a significantly broader class of systems, referred to as generically tilted multi-rotors, admitting an arbitrary number of actuators and diverse geometric configurations. The objective of this line of work is to enable static hovering even in the presence of propeller failures, including configurations in which translational and rotational dynamics are intrinsically coupled. The proposed approach departs from the classical assumption that the control force is orthogonal to the plane containing the propeller centers and that control moments can be generated independently of the thrust force. Instead, the authors consider the more general case in which the direction of the control force is not fixed and control moments may induce translational effects. Within this framework, they introduce the concept of zero-moment direction (also referred to as preferential direction), defined as a virtual direction along which the intensity of the control force can be freely assigned while maintaining a zero net control moment. While this formulation provides a potentially unifying control paradigm for a wide range of multirotor configurations, its stability properties are supported primarily by heuristic arguments. Nevertheless, extensive experimental validation has demonstrated a remarkably large basin of attraction in fault scenarios. In particular, the controller was shown to successfully recover stability despite transient attitude errors approaching $50^{\circ}$, angular rates exceeding $50^{\circ}/\mathrm{s}$, and velocity errors above $1\,\mathrm{m/s}$ [15].
A subsequent work [13] addresses a similar problem by providing sufficient conditions for local asymptotic stability of a hovering equilibrium within the same general modeling framework. However, the analysis introduces additional feasibility requirements on the actuation capabilities that effectively restrict the result to configurations in which the translational component induced by moment allocation can be locally eliminated by an equivalent input parametrization. Therefore, the stability guarantees of [13] do not apply to generic platforms with spurious forces.
Therefore, despite the strong empirical evidence provided in [15], and the theoretical contribution provided in [13], the stability analysis of a controller applicable to the class of floating rigid bodies with spurious forces (such as the one presented in [14]) is still an open problem.
@@
Our contribution provides, for the first time, a rigorous Lyapunov-based stability proof for the geometric control law proposed by [14], extending its validity to the abstract canonical model presented in Section II. Our analysis relaxes the feasibility requirements [13], thereby being applicable to the class of floating vehicles with spurious forces. This result closes the gap between prior empirical evidence [14] and theoretical certification. We prove that the controller guarantees local exponential stability of the hovering equilibrium. This work focuses exclusively on the theoretical certification, serving as the formal counterpart to the experimental studies of [14].
@@
From the methodological point of view, the proposed analysis is a non-trivial extension of the methodology presented in Lee et al. [8, 9]. The presence of spurious forces fundamentally alters the system structure compared to [8]. In particular, the induced non-minimum-phase behavior and intrinsic coupling violate the skew-symmetry and cascade properties on which the analysis of [8] relies. To overcome these structural obstacles, we introduce a reformulated Lyapunov argument with new bounds specifically designed to handle the additional coupling terms.
@@
The remainder of the work is organized as follows. Section II introduces the generic canonical model and summarizes the control law. Section III details the stability analysis. Concluding remarks are provided in Section IV and the Appendix A collects some of the identities used throughout the work.